OpenStudy (anonymous):

Solve each radical equation by using algebra. Check your solution. I will use sq. rt for square root. 1. sq. rt of (x- 5) = 3 2. sq. rt of ( 2x - 1) = x 3. 3sq. rt of( x - 2) = 2sq. rt of( x) 4. sq. rt. of (3x - 2) = x - 2

OpenStudy (unklerhaukus):

\[\sqrt{x-5}=3\]\[ x-5=3^2\]\[x=\cdots\]

OpenStudy (lgbasallote):

@GoDeanoGo do you know how to solve that ^

OpenStudy (anonymous):